OpenStudy (anonymous):

any idea how to get the zeros for x^3-4x^2+x+6 without using a graphing calculator? (can't figure out how to factor it out)

OpenStudy (anonymous):

I should say algebraically, I know I can make a table and just get lucky

OpenStudy (dumbcow):

try to factor out (x+-1), (x+-2)...up to 6 if that works you'll be left with a quadratic

OpenStudy (dumbcow):

actually you can factor out (x+1) there are different ways of doing this (x+1)(x^2 +bx +6) = x^3-4x^2+x+6 we have to find b, if b is an integer it works x^2+bx^2 = -4x^2 (b+1)x^2 = -4x^2 b+1 = -4 b = -5

OpenStudy (anonymous):

ah long division, its been a while

OpenStudy (dumbcow):

long division always works too :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You could also use the rational root theorem to discover the possible rational roots.